# Offline Reinforcement Learning with Imbalanced Datasets

### v1

@@ 3.1 State-Agnostic Pessimism
Refer to caption Refer to caption Refer to caption
@@ 3.2 Uniformly Sampling
Refer to caption Refer to caption Refer to caption
@@ 6 Related Work
Offline RL without dataset consideration. To alleviate the above problem, previous methods enforce pessimism in actor updates (Wu et al., 2019; Fujimoto et al., 2019; Kumar et al., 2019a; Fujimoto & Gu, 2021; Xu et al., 2021; Zhou et al., 2021) or critic updates (Liu et al., 2020; Kumar et al., 2020; Bai et al., 2022), whose learned policy would query the Q-functions of out-of-distribution actions to outperform the behavior policy, implicitly or explicitly minimizing the divergence over those two policies to mitigate the distributional shift problem. Another branch of methods learns an optimal policy by using only in-sample transitions, also called in-sample learning, bypassing querying the values of unseen actions. This can be achieved by vanilla advantage-weighted regression (Peng et al., 2019; Nair et al., 2020), the extreme approximation of advantage-weighted regression (Chen et al., 2020; Kostrikov et al., 2021; Brandfonbrener et al., 2021), conditioning on some future information or target information (Kumar et al., 2019b; Chen et al., 2021; Emmons et al., 2021), or find the best next states in datasets (Xu et al., 2022). Those methods have achieved significant success in standard benchmarks (Gulcehre et al., 2020; Fu et al., 2020; Qin et al., 2021), however, they have not adequately considered the properties of the dataset, despite it being the primary source to extract policies. In this study, we shift our focus towards offline RL for real-world applications by taking into account the distribution of real-world datasets (i.e., imbalanced datasets), and examining the limitations of current algorithms in such scenarios.

### v3

@@ 6 Related Work
Offline RL without dataset consideration. To alleviate the above problem, previous methods enforce pessimism in actor updates (Wu et al., 2019; Fujimoto et al., 2019; Kumar et al., 2019a; Fujimoto & Gu, 2021; Xu et al., 2021; Zhou et al., 2021) or critic updates (Liu et al., 2020; Kumar et al., 2020; Bai et al., 2022; Li et al., 2022), whose learned policy would query the Q-functions of out-of-distribution actions to outperform the behavior policy, implicitly or explicitly minimizing the divergence over those two policies to mitigate the distributional shift problem. Another branch of methods learns an optimal policy by using only in-sample transitions, also called in-sample learning, bypassing querying the values of unseen actions. This can be achieved by vanilla advantage-weighted regression (Peng et al., 2019; Nair et al., 2020), the extreme approximation of advantage-weighted regression (Chen et al., 2020; Kostrikov et al., 2021; Xu et al., 2023; Brandfonbrener et al., 2021), conditioning on some future information or target information (Kumar et al., 2019b; Chen et al., 2021; Emmons et al., 2021; Feng et al., 2022), or find the best next states in datasets (Xu et al., 2022). Those methods have achieved significant success in standard benchmarks (Gulcehre et al., 2020; Fu et al., 2020; Qin et al., 2021), however, they have not adequately considered the properties of the dataset, despite it being the primary source to extract policies. In this study, we shift our focus towards offline RL for real-world applications by taking into account the distribution of real-world datasets (i.e., imbalanced datasets), and examining the limitations of current algorithms in such scenarios.